# Online Multivariate Changepoint Detection: Leveraging Links With Computational Geometry

### v2

@@ 4.4.1 MdFOCuS(0) dyadic: a more complex yet provably quasi-linear online algorithm
We empirically study the complexity of Algorithm 3 in Figure 16 of Supplementary Material F. It can be seen that it is slower than Algorithm 1. Algorithm 3 considers at least all points on the hull to maximise the likelihood for every new data point. For a fixed $p$ following, Theorem 6 and its corollary the number of points on the hull at step $i$ is in $\Omega((\log(i))^{p}).$ To recover a lower bound on the complexity, we first sum overall $i\leq n$, and then consider only $i\geq n/2$ to recover that the overall number of points we will consider is in $\Omega(n/2(\log(n/2))^{p})$ and thus $\Omega(n(\log(n))^{p}).$ Relative to that Algorithm 3 loose just one $\log(n)$ term, we thus argue our bound is relatively tight.
4.4.2 MdFOCuS(0) Sp: A simple online algorithm with sparse statistics
Out of simplicity in lines 9 and 10 of Algorithm 1 we only consider the full (or dense) max-likelihood ratio statistics, as in Equation (6), that is assuming all coordinates are changing. However, based on Theorem 5 the changepoints maximising sparse likelihood ratio statistics, that is assuming only a subset of $s$ coordinates are changing (see Equation (2)), also correspond to points on the hull.
Computing all sparse max-likelihood ratio statistics, as in Equation (2), is a bit more difficult than the full likelihood ratio statistics. To be specific, for each candidate changepoint we sort the per-dimension ratio statistics (this can be done in $\mathcal{O}(p\log(p))$) and we then sum these from the largest to the smallest in $\mathcal{O}(p)$ to iteratively recover the best for all sparsity level $s.$ Once this is done for all candidates, we recover the maximum for each sparsity level in $\mathcal{O}(\mathcal{T}).$ Adding these up, even if we consider all sparsity levels (that is all $s$ between 1 and $p$) we are overall in $\mathcal{O}(|\mathcal{T}|p\log(p)).$
In practice, it makes sense to consider a subset of all sparsity levels. In our simulations we consider: $s\in\{1,2,\ldots,2^{\lfloor\log_{2}(p)\rfloor},p\}.$ In the spirit of our FOCuS approximation [34] we also consider the sum of the max statistics: that is we get for each coordinate the max likelihood ratio statistics and sum these. We call these procedures respectively MdFOCuS0 Sp if we know the pre-change parameter and MdFOCuS Sp if not. As can be seen in Figure 5 (left), the empirical runtime of MdFOCuS Sp is slower than that of MdFOCuS and comparable to that of the ocd R package [38].
4.4.3 MdFOCuS(0) $\tilde{p}$d Sp: An approximate algorithm for large $p$
Under the condition of Theorem 6, Algorithm 1 is expected to store $\mathcal{O}(\log^{p}(n))$ candidates and the $\log^{p}(n)$ terms means that the algorithm is impractical for $p$ larger than $5,$ as exemplified in Figure 4. In Supplementary Material G we propose an approximate algorithm by recovering only a subset of the points on the hull. We obtain this subset by considering projections on at most $\tilde{p}$ variates, running the hull algorithm on each projection, and taking the union of the time-points (of the vertices) retained for each projection. For each projection or group of variates, and under the condition of Theorem 6 we thus store $\mathcal{O}(\log^{\tilde{p}}(n))$ candidates. Considering $p$ such groups we thus store $\mathcal{O}(p\log^{\tilde{p}}(n))$ candidates in total thus effectively controlling the quality of the approximation and reducing the computational complexity.
This heuristic for $\tilde{p}=2$, implemented in R, runs on average in about 100 seconds for $p=100$ and $n=10000$. We further explore the run times of this heuristic in Figure 5 and show that its run time is comparable to ocd [38]. Recovering just $\mathcal{O}(p\log^{2}(n))$ points of the hull out $\mathcal{O}(p\log^{{p}}(n))$ is arguably a crude approximation, but as shown in Section 4.6 it already provides good statistical performance relative to ocd or the FOCuS(0) multivariate approximation.
@@ 4.4.3 MdFOCuS(0) $\tilde{p}$d Sp: An approximate algorithm for large $p$
4.5 A Note on Statistical Theoretical Guarantees
In Corollary 2, in Supplementary Material H, we provide time-varying thresholds (i.e. increasing with the number of observations $n$) that control the probability of incorrectly detecting a change (i.e. probability of false alarms) for the i.i.d. Gaussian setting. These results are for considering only a sparse max likelihood ratio statistics with $s=1$ and the full max likelihood ratio statistics (that is with $s=p$). We also provide fixed thresholds to control the average run length. Further, in Lemma 3 we derive a bound on the average detection delay (ADD) for a fixed threshold.
We derive the thresholds following the peeling approach of [65] combined with the Chi-square concentration bound of [66]. For the sparse statistics our threshold is $8\log(np/\alpha)$ and for the full or dense statistics it is $p+2\sqrt{p(4\log(n)-\log(\alpha)+1)}+4\log(n)-\log(\alpha)+1,$ where $2\alpha$ is the desired false positive detection rate. These thresholds are comparable to the $T^{diag}$ and $T^{off}$ thresholds of ocd [38].
Note that, in Supplementary Material, we only consider two sparsity levels: $s=1$ and $s=p.$ In MdFOCuS Sp we typically consider more, say $s$ in $1,2,\dots,2^{\lfloor\log_{2}(p)\rfloor}$. Here is one way to extend the thresholds to this case. First, for any sparsity $s$ strictly larger than 1 and smaller than $p$, we consider a union bound over all possible choices of $s$ coordinates among $p$, which is (${p\choose s}$ (see the partial norm statistics in [67]). Second, we consider a union bound over all $\log_{2}(p)$ sparsity levels. Combining these, we recover a $\log(\log_{2}(p))$ term in all thresholds (as in the $T^{\text{diag}}$ and $T^{\text{off}}$ thresholds of ocd) and a $\log({p\choose s})\leq s\log(ep/s)$ term in thresholds for sparsity $s$ strictly larger than $1$ and smaller than $p.$
For the ADD our bound is inversely proportional to the size of the change. Combined with the threshold, this supports the intuition that if a change is in just one dimension it should be detected by a partial statistics ($s=1$) while if it is dense it should be detected by full maximum likelihood ratio statistics ($s=p$). In the dense case, our bound on the ADD is proportional to the threshold and therefore includes a term linear in $p.$ This is likely sub-optimal and due to our proof technique based on a martingale argument. We show in Subsection H.2.1 in Supplementary Material exploiting the peeling argument of [65] and [68] that with probability at least $1-\alpha$ the detection delay is smaller than a constant only involving a $\sqrt{p}$ terms divided by the size of the change.
4.6 Empirical statistical performances
We evaluate the empirical performance of the R implementation of MdFOCuS Sp (which computes the sparse likelihood ratio statistics) on synthetic data (the link to the code is available in the Supplementary Material). In this section, we analyse 3, 5, and 100-dimensional time series under the Gaussian change-in-mean setting, with an additional simulation study for 50-dimensional time series presented in Supplementary Material E.3. We consider two scenarios: pre-change mean known and unknown. For each experiment, we report results aggregated from 300 data sets.
For MdFOCuS Sp for dimension $p=3$ and $5$, we consider sparsity level $s=1,$ the full max-likelihood ratio (either $s=3$ or $s=5$ depending on the dimension), and the sum of the max likelihood ratio statistics per dimension as proposed in [34] (see Subsection 4.4). For higher dimensions ($p=100$ and $50$), we set the sparsity levels as in Section 4.4. We compare the performance of MdFOCuS with the high-dimensional online changepoint detection method ocd from [38] and a naive multivariate implementation of FOCuS, obtained by running FOCuS independently for each dimension and summing or taking the maximum of the resulting traces (see Section 6 of [34]) at each iteration. Additionally, for the pre-change mean unknown case, we report results of FOCuS and ocd where a plug-in estimate of the pre-change parameter was obtained over a probation (training) period of 250 observations.
@@ 4.6 Empirical statistical performances
Thresholds were selected under a Monte-Carlo simulation to achieve an average run length of $5\times 10^{3}$ observations with no change, consistent with the empirical evaluation studies of both [34] and [38]. An additional study controlling the false positive rate can be found in Supplementary Material E.3. For a change at $t=1000$, we evaluate performances for a range of different change magnitudes and density regimes, e.g., for a change affecting only a portion of the time series. To compare performances across different density regimes, changes in the affected time series were normalised such that the change direction lies in the $p$-dimensional sphere, i.e., ($||\eta_{1}-\eta_{2}||^{2}_{2}=\delta$), where $\delta$ is the overall change magnitude and $\eta_{1},\eta_{2}$ are the pre- and post-change means, respectively.
We report below tables summarising the ADD over various magnitudes and for different change densities across all replicates, respectively for 3-dimensional time series (in Table 2), 5-dimensional time series (in Table 3), and finally 100-dimensional time series (in Table 4).
In all tables, the two leftmost columns indicate the scenario in terms of change magnitude and density. In the pre-change parameter known scenario, on the left, with FOCuS0, MdFOCuS0 Sp, and ocd (oracle), we indicate the respective algorithms with plug-in oracle pre-changepoint parameter. On the right, for the pre-change mean unknown scenario, with FOCuS0 (est 250) and ocd (est 250), we report the pre-change mean known methods, using a plug-in pre-change mean estimate, and with FOCuS, the independent FOCuS algorithm, and finally MdFOCuS Sp, the procedure introduced in this manuscript. Finally, for the 100-dimensional profiles, with MdFOCuS0 2d Sp and MdFOCuS 2d Sp, we indicate that the procedure is employing the 2d approximation for the convex hull.
@@
Globally, for all dimensions, in both pre-change parameter known and unknown scenarios, we observe that MdFOCuS Sp and MdFOCuS0 Sp generally achieve smaller detection delays in cases where the change affects all the time series and over changes of larger magnitudes. In scenarios with density 1, the simple FOCuS aggregation tends to be slightly better. This is expected as if we only have a change in one dimension, both the sparse statistics of FOCuS and MdFOCuS Sp will be the same, but MdFOCuS Sp will have a slightly larger threshold, resulting in a slower delay.
@@
Focusing on the pre-change mean known case, over changes of smaller magnitudes, we find MdFOCuS0 Sp to be relatively close to the best competitor in terms of ADD, while it generally outperforms the other methods over changes of larger magnitudes. Furthermore, MdFOCuS0 Sp is seen to have quite consistent performances across different sparsity regimes over the same change magnitudes. As ocd uses a grid for the possible size of change for each variate, its performance is affected by how close the true change size is to the nearest point on the grid. In particular, its performance deteriorates when the true change is outside the set of grid values. We observe this in its deterioration of performance for low magnitudes (0.125) and large magnitudes (1 or 2), in particular over dimensions $p=3,5$. This effect emerging from a grid-based approximation has already been observed thoroughly, albeit in the 1-d case, in the main simulation study of [34].
@@
Moving to the pre-change mean unknown scenario, we observe that MdFOCuS Sp generally outperforms competitor methods for magnitudes larger than 0.250 in most scenarios, with the exception of scenarios with a change in a single dimension, as explained above. Moreover, in dimensions $p=5$ and $100$, we can clearly see the advantages of performing the likelihood ratio test for a change when the pre-change mean is unknown over FOCuS0 or ocd, where the pre-change mean unknown parameter needs to be estimated. The only exception occurs in the case of FOCuS (est 250) for $p=3$, where in the lowest magnitude, it is seen outperforming all other methods. Whilst this might seem counter-intuitive, as the error from the estimation should propagate, in dimension $p=3$ for a small signal-to-noise ratio, the estimation is good enough to not impair the average run length and increase the thresholds, while benefiting from an easier optimisation (as we see in the pre-change mean known case). However, as we move to higher dimensions $p=5$ and $100$, it is more unlikely to obtain a better estimate of the true pre-change parameter simultaneously in all dimensions. This results in larger thresholds to control the desired run length, and therefore in a global loss of power for FOCuS0 (est) and ocd (est). Of course, having more training observations improves the results, in particular for $p=3$, however, this still has little to no effect in higher dimensions (we report results with more training observations in Supplementary Material E.4). Moreover, in Supplementary Material E.3, where we control for false positives and have a much later change, it can be seen in Table 5, that our 2d approximation MdFOCuS 2d Sp is actually better than the 1d approximation of FOCuS [34] for dense changes showing the advantage of considering larger $\tilde{p}.$
@@
Table 2: Average detection delay with 3-dimensional time series for a change at time $t=1000$. Pre-change parameter is known on the left, and unknown on the right. Best results per row are highlighted in bold.
@@
magnitude density FOCuS0 MdFOCuS0 Sp ocd (ora) FOCuS0 (est 250) FOCuS MdFOCuS Sp ocd (est 250)
@@
Table 3: Average detection delay with 5-dimensional time series for a change at time $t=1000$. Pre-change parameter is known on the left, and unknown on the right. Best results per row are highlighted in bold.
@@
magnitude density FOCuS0 MdFOCuS0 Sp ocd (ora) FOCuS0 (est 250) FOCuS MdFOCuS Sp ocd (est 250)
@@
Table 4: Average detection delay with 100-dimensional time series for a change at time $t=1000$. Pre-change parameter is known on the left, and unknown on the right. Best results per row are highlighted in bold.
@@
magnitude density FOCuS0 MdFOCuS0 2d Sp ocd (ora) FOCuS0 (est 250) FOCuS MdFOCuS 2d Sp ocd (est 250)
@@
5 Application on NBA Plus-Minus
In this section, we illustrate our methodology to detect a change in the Cleveland Cavaliers (an NBA team) Plus-Minus score from season 2010-11 to 2017-18. This example was proposed in [37]. We use this example for two reasons. First, it highlights that our computational framework is also applicable to CUSUM Exponential baseline e-detectors proposed in [37] which in essence are a non parametric generalization of generalized likelihood ratio rules. Second, it illustrates how modelling not only the signal’s mean but also its variance can simplify the calibration of the detection threshold, without necessarily resorting to a non-parametric approach.
We recovered the Plus-Minus score using the nbastatR package [69] available on GitHub https://github.com/abresler/nbastatR. The Cavaliers Plus-Minus is represented in Figure 6. This score represents how well the Cavaliers fared against their opponent. In this subsection, we assume that we are observing these matches one after the other and we are interested in detecting a changepoint online.
@@ 5 Application on NBA Plus-Minus
As detailed in [37] it is nontrivial to fit this problem into commonly used parametric sequential changepoint detection procedures for two main reasons. First, it is not easy to choose a parametric model to fit the score. In particular, we would like to take into account that we have integer-valued data, and we expect some form of dependency. Second, assuming we have chosen a model or likelihood function, calibrating the threshold to detect a changepoint is not trivial. A small change in the distribution we use as a reference can substantially change the threshold and as a consequence increase the detection time or lead to false detections. E-detectors of [37], are an interesting non-parametric solution to these two problems assuming some knowledge of the pre-change distribution.

### v3

@@ 4.5 Two variant of MdFOCuS
4.6 Statistical Guarantees
In Supplementary Material I, we provide several statistical guarantees for our procedure considering sparse or dense statistics in the multivariate Gaussian case. Our results also cover test statistics that threshold the contribution from each component. We consider properties of our algorithm if there is no change and then if there is one. In the first scenario, we cover both the pre-change parameter being known or unknown, while in the second, we only consider the pre-change parameter being known. We summarise the main results and their implications here.
We first consider the properties of our algorithm if there is no change. In this case, one may want to control either the false-positive rate, that is, the probability of ever detecting a change, or the average run length, the expected time until we detect a change. Our results give thresholds for both cases, but we will summarise them here just for an average run length of $\gamma$. At time $n$, we define the test statistic assuming a sparsity level $s$ and a candidate change at $\tau$ as $S^{s}_{\tau,n}$. Testing for $m$ different sparsity levels, and with the following threshold for sparsity $s$
$c_{s}(\gamma)=s+2\zeta+2\sqrt{s\zeta},\mbox{ where }\zeta=4\log\gamma+s\log p+\log m+5\log 2$
our MdFOCus algorithm will have an average run length of at least $\gamma$ (this follows by applying the argument in Remark 4 to the result of Proposition 8 with $N=2\gamma$ and $\alpha=1/2$). In the case where $s=p$ the $s\log(p)$ term in $\zeta$ can be omitted.
Next, we consider the power of detecting a change. This can be quantified in terms of the average detection delay, informally the expected time after a change that you would detect it, or bounds on the detection delay, the probability of detecting a change within some time-interval that it occurs. We provide results on both, but will summarise here only the latter. Let the change in mean be specified by the vector $\delta$. For a test at sparsity level $s$, denote by $\delta^{[s]}$ the vector of length $s$ that consists of the $s$ entries of $\delta$ with the largest absolute value. Finally, let $\tau^{*}$ denote the true changepoint and $T^{s}$ the time at which we detect a change using the $s$-sparse statistic. Then from Proposition 9 and Remark 7, and using the threshold $c_{s}(\gamma)$, we have that with probability at least $(1-\alpha)$,
$T^{s}-\tau^{*}<\frac{2(c_{s}(\gamma)-s)+2\sqrt{s\log(3/\alpha)}-8\log(\alpha/3)}{\|\delta^{[s]}\|^{2}}.$ (11)
We can draw some insights from these results. Excluding $s=p$, first observe that $c_{s}(\gamma)$ is increasing with $s$. This means, unsurprisingly, that to minimise the bound on the detection delay $T^{s}-\tau^{*}$, we never want to use $s$ greater than the true sparsity of $\delta$, with the notable exception of $s=p$. More generally, it is helpful to consider two different regimes depending on the relative sizes of $\gamma$ and $p$. The first is a low-dimensional setting where $\log\gamma$ is much larger than $p$. In this case, the numerator of the right-hand side of (11) is dominated by the $\log\gamma$ term in $c_{s}(\gamma)$, and will be approximately $8\log\gamma/\|\delta^{[s]}\|^{2}$. In this case, maximising $\|\delta^{[s]}\|^{2}$ is key to minimising the detection delay, and this can be guaranteed by using the dense test statistic, $s=p$. Any advantage of using smaller $s$, in the case that some components do not change, will be small.
The other regime is where $p$ is much larger than $\log\gamma$. In this case, using the dense statistics will give a detection delay with leading order (from the right-hand side of Equation 11) that is proportional to $\sqrt{p}/\|\delta\|^{2}$. By comparison, the $s$-sparse statistic will have a detection delay with leading order proportional to $(s\log p)/\|\delta^{[s]}\|^{2}$. If only $s$ components change, and all change by the same amount, then this latter quantity will be lower order in $p$ when $s=o(\sqrt{p})$, as is consistent with other results on high-dimensional change detection (e.g. Chen et al., 2022). Finally, we can see that the benefit of robustness of performance to the true level of sparsity that comes using $m=O(\log p)$ tests for differing sparsity levels is only a $\log\log p$ increase in the thresholds and hence in the detection delays.
4.7 Empirical statistical performance
In this section, we evaluate the empirical performance of the R implementation of MdFOCuS with the sparse statistics (as in Section 4.5) on synthetic 3, 5, and 100-dimensional time series under the Gaussian change-in-mean setting, with an additional simulation study for a 50-dimensional time series presented in Supplementary Material F.3. We consider two scenarios: pre-change mean known and unknown. For each experiment, we report results aggregated from 300 data sets.
@@ 4.7 Empirical statistical performance
Thresholds were selected under a Monte-Carlo simulation to achieve an average run length of $5000$ observations. An additional study controlling the false positive rate can be found in Supplementary Material F.3. For a change at $t=1000$, we evaluate performances for a range of different change magnitudes and sparsity levels. To compare performances across different sparsity regimes, changes in the affected time series were normalised such that the change direction lies in the $p$-dimensional sphere, i.e., ($||\eta_{1}-\eta_{2}||^{2}_{2}=\delta$), where $\delta$ is the overall change magnitude and $\eta_{1},\eta_{2}$ are the pre- and post-change means, respectively. Tables 1 to Table 3 show results of the Average Detection Delays (ADDs) for 3, 5, and 100-dimensional time series, respectively.
@@
For all dimensions of time series, in both pre-change parameter known and unknown scenarios, we observe that MdFOCuS generally achieves smaller detection delays in cases where the change affects all the time series and over changes of larger magnitudes. In scenarios with sparsity set to 1, the simple FOCuS aggregation tends to be slightly better. This is expected as if we only have a change in one dimension, both the sparse statistics of FOCuS and MdFOCuS will be the same, but MdFOCuS will have a slightly larger threshold, resulting in a slower delay.
@@
For the pre-change mean known case, over changes of smaller magnitudes, we find MdFOCuS to be relatively close to the best competitor in terms of ADD, while it generally outperforms the other methods for larger changes. Furthermore, MdFOCuS is seen to have consistent performances across different sparsity levels. As ocd uses a grid for the possible size of change for each variate, its performance is affected by how close the true change size is to the nearest point on the grid. In particular, its performance deteriorates when the true change is outside the set of grid values. We observe this in its deterioration of performance for low magnitudes (0.125) and large magnitudes (1 or 2), in particular over dimensions $p=3,5$.
@@
For the pre-change mean unknown scenario, we observe that MdFOCuS generally outperforms competitor methods for magnitudes larger than 0.250, with the exception of scenarios with a change in a single dimension, as explained above. Moreover, in dimensions $p=5$ and $100$, we can clearly see the advantages of performing the likelihood ratio test for a change when the pre-change mean is unknown over using a plug-in estimate for the pre-change parameter. The only exception occurs in the case of FOCuS, when estimating the pre-change parameter from the training data, for $p=3$, which out-performs other methods for the smallest size of change. Whilst this might seem counter-intuitive, as the error from the estimation should propagate, in dimension $p=3$ for a small signal-to-noise ratio, the estimation is good enough to not impair the average run length and increase the thresholds, while benefiting from an easier optimisation (as we see in the pre-change mean known case). However, as we move to higher dimensions $p=5$ and $100$, it is unlikely to obtain a better estimate of the true pre-change parameter simultaneously in all dimensions. This results in larger thresholds to control the desired run length, and therefore in a global loss of power. Of course, having more training observations improves the results, in particular for $p=3$, however, this still has little to no effect in higher dimensions (see Supplementary Material F.4).
@@
Table 1: Average detection delay with 3-dimensional time series for a change at time $t=1000$. Pre-change parameter is known on the left, and unknown on the right. Best results per row are highlighted in bold. For the pre-change unknown case FOCuS (est 250) and ocd estimate the pre-change parameter from 250 training samples.
@@
magnitude sparsity FOCuS MdFOCuS ocd FOCuS (est 250) FOCuS MdFOCuS ocd (est 250)
@@
Table 2: Average detection delay with 5-dimensional time series for a change at time $t=1000$. Pre-change parameter is known on the left, and unknown on the right. Best results per row are highlighted in bold. For the pre-change unknown case FOCuS (est 250) and ocd estimate the pre-change parameter from 250 training samples.
@@
magnitude sparsity FOCuS MdFOCuS ocd FOCuS (est 250) FOCuS MdFOCuS ocd (est 250)
@@
Table 3: Average detection delay with 100-dimensional time series for a change at time $t=1000$. Pre-change parameter is known on the left, and unknown on the right. Best results per row are highlighted in bold. For the pre-change unknown case FOCuS (est 250) and ocd estimate the pre-change parameter from 250 training samples.
@@
magnitude sparsity FOCuS MdFOCuS ocd FOCuS (est 250) FOCuS MdFOCuS ocd (est 250)
@@
5 Application to NBA Plus-Minus Data
In this section, we illustrate our methodology to detect a change in the Cleveland Cavaliers (an NBA team) Plus-Minus score from season 2010-11 to 2017-18. This example was proposed in Shin et al. (2022). We use this example for two reasons. First, it highlights that our computational framework is also applicable to CUSUM Exponential baseline e-detectors proposed in Shin et al. (2022) which in essence are a non parametric generalization of generalized likelihood ratio rules. Second, it illustrates how modelling not only the signal’s mean but also its variance can simplify the calibration of the detection threshold, without necessarily resorting to a non-parametric approach.
We recovered the Plus-Minus score using the nbastatR package Bresler (2023) available at https://github.com/abresler/nbastatR. The Cavaliers Plus-Minus is represented in Figure 5, and is a time series of the difference in score of the Cavaliers and their opponents over successive games. We consider detecting a changepoint in this data online.
@@ 5 Application to NBA Plus-Minus Data
As detailed in Shin et al. (2022), it is nontrivial to fit this problem into commonly used parametric sequential changepoint detection procedures for two main reasons. First, it is not easy to choose a parametric model to fit the score. In particular, we would like to take into account that we have integer-valued data, and we expect some form of dependency. Second, assuming we have chosen a model or likelihood function, calibrating the threshold to detect a changepoint is not trivial. A small change in the distribution we use as a reference can substantially change the threshold and as a consequence increase the detection time or lead to false detections. E-detectors Shin et al. (2022), are a non-parametric solution to these two problems.